Find  f −1(x). f(x) = 4x − 8 f −1(x) =

Find  f −1(x).

f(x) = 4x − 8

f −1(x) =
Answer #1

Given fo) = 42-8 let fogy= 42-8 y = 428 y+8 = 48 y+8 = x 4 4+2 f (2) is obtained by replacing y with x fo)= 4x+2